5 ul (10ng.ul) of ptud plasmid control, 2 colonies grown

100ng of DNA was used for ligation, 25ng DNA was used for transformation, 1ml was the total volume of cells grown, 100 ul was plated, 130 colonies were grown

1. What was your transformation efficiency for control DNA pTUD at concentration of 10 ng/ul)?

2. What was transformation efficiency for experimental ligation (you will need to calculate approximate concentration from the steps of the cloning procedure you performed)?
Answer #1

1) Transformation efficiency for control DNA pTUD at concentration of 10ng/ul = number of colonies/amount of dna= 2/50 = 0.04 cfu/ng = 40 cfu/ug

2) Transformation efficiency for experimental ligation:

Amount of dna = vol spread on plate * dna used / total volume of transformation = 100*25/1000= 2.5 ng

Transformation efficiency = colonies/dna used = 130/2.5 = 52 cfu/ng= 52000 cfu/ug